OpenStudy (anonymous):

Using complete sentences, explain how solving a literal equation is similar to or different from simplifying an expression such as 6 - 2(52 + 7) ÷ 4.

OpenStudy (anonymous):

math teacher problem

OpenStudy (anonymous):

a literal equation is an equation, such as Area of a Square = Length x Height, \[A=L*H\]

OpenStudy (anonymous):

for instance, to solve that literal equation for height, if you are given Area and Length, you would have \[H = A/L\]

OpenStudy (anonymous):

the expression 6 - 2*(52+7) / 4 has no variables in it and it simply equals a value

OpenStudy (anonymous):

\[6-(2*59)/4\] = \[6-29.5\] = \[-23.5\]

OpenStudy (anonymous):

they are similar in that you still follow the order of operations

OpenStudy (anonymous):

Groupings, Exponents, Multiplication, Division, Addition

OpenStudy (anonymous):

so a literal equation contains variables